OpenStudy (anonymous):

Which equation represents the line passing through the points (1, 1) and (−2, 13)? 4x − y = 5 x − 4y = 5 x + 4y = 5 4x + y = 5

OpenStudy (anonymous):

@shkrina can you helpp?

OpenStudy (anonymous):

I dont really understand this, so can you just tell me the answer please?

OpenStudy (anonymous):

use the eq formula ...

OpenStudy (anonymous):

y -y1 x-x1 -------- = ---------- y2-y1 x2-x1 where (x1,y1)=(1,1) (x2, y2)=(-2,13)

Directrix (directrix):

@christinechloe21 There are two option lines on which the point (1, 1) lies. Look at the third option: x + 4y = 5 Substitute 1 for x and 1 for y and see if what results is a true statement. 1 + 4*1 = 5 1 + 4 = 5 5=5 On the fourth option, do the same and you will get 5= 5. ------------ Now follow the same procedure for the point (-2,13) and first the line with equation: x + 4y = 5 -------------- If that works, that is, if -2 + 4*13 = 5, then x + 4y=5 is the answer. ---------------- If not, the test the last option line 4x + y = 5 to determine if (-2, 13) makes that equation a true statement when you substitute -2 for x and 13 for y. If the resulting statement is true, that is the answer to the question.